On December 30, 2020, Sherlock Co. acquired 100% of Barrett Corporation for $750,000 cash. The post-combination, but pre-consolidation, balance sheets of the two firms showed total assets of $1,500,000 for the parent and $440,000 for the subsidiary. The book values of the subsidiary’s net assets approximated their fair values, and there are no unrecorded net assets.

Total assets on the consolidated balance sheet would be:

a. $   750,000

b. $1,190,000

c. $1,940,000

d. $2,690,000

On January 1, 2020, Sarai, Inc. purchased 40% of the voting common stock of Rolls Corp. for $500,000. There was no amortization. During 2020, Rolls paid dividends of $50,000 and reported a net loss of $120,000.

What is the balance in the Equity Investment account on December 31, 2020?

a. $432,000

b. $380,000

c. $528,000

d. $548,000

Spring Creek, Inc. purchased a 18% interest in Floyd Corporation on January 2, 2020. The purchase price was $200,000. Spring Creek's officers constitute a majority of Floyd Corporation’s board of directors. The investee reported net income of $300,000 and paid dividends of $50,000 in 2020.

On the December 31, 2020, balance sheet, what amount should Spring Creek, Inc. report as Equity Investment in Floyd Corporation?

a. $245,000

b. $254,000

c. $263,000

d. $300,000
Answer #1

Sherlock Co.

Answer: b. $1,190,000

Total assets on the consolidated balance sheet = $1,500,000 + $440,000 - $750,000 = $1,190,000

Sarai Inc.

Answer: a. $432,000

Balance in Equity Investment account on December 31, 2020 = $500,000 – (40% x $50,000) - (40% x $120,000) = $500,000 – 20,000 - $48,000 = $432,000

Spring Creek, Inc.

Answer: a. $245,000

Since Spring Creek’s officers constitute a majority of Floyd Corporation’s board of directors, the circumstances indicate Spring Creek to have significant influence over Floyd Corporation and hence the investment in Floyd Corporation will be accounted for under the equity method.

Equity Investment in Floyd Corporation on December 31, 2020 balance sheet = $200,000 – (18% x $50,000) + (18% x $300,000) = $200,000 - $9,000 + $54,000 = $245,000.